OpenStudy (anonymous):

The point is on the terminal side of an angle in standard position. Determine the exact values of the six trigonometric values of the angle. (2/3, 5/8)

OpenStudy (anonymous):

I know how to get the sin, cos, tan and their reciprocals, but I'm having trouble with the fractions. I know that you need to apply \[r=\sqrt{x^2+y^2}\] first, but getting the answer is stumping me.

OpenStudy (aum):



OpenStudy (aum):

\[r^2 = \left (\frac 23\right )^2 + \left(\frac 58 \right)^2 = \frac {4}{9} + \frac{25}{64} \]

OpenStudy (anonymous):

And this is where I am stuck

OpenStudy (aum):

\[r^2 = \frac {4}{9} + \frac{25}{64} = \frac{4*64+25*9}{9*64} = \frac{256+225}{9*64} = \frac{481}{9*64} \\ r = \frac{\sqrt{481}}{3*8}= \frac{\sqrt{481}}{24}\]

OpenStudy (anonymous):

ah, fractions. always hated these. can you explain to me what exactly happened there?

OpenStudy (aum):

\[\frac ab + \frac cd = \frac{ad + bc}{bd}\]

OpenStudy (anonymous):

and here I am, back to algebra 1. totally forgot about the identities of fractions. so now we have \[\frac{ \sqrt{481} }{ 24 }\] then what? would we just apply sin= y/r, cos=x/r and tan = y/x and their inverses?

OpenStudy (aum):

Exactly. And if you get a radical in the denominator, get rid of it by multiplying top and bottom by the radical in the denominator.

OpenStudy (aum):

Many teachers and textbooks don't like radicals in the denominator and so you have o rationalize the denominator.

OpenStudy (bossimbacon):

@aum can you plz help me I'm sorry @bbbbbrrrruuuuhhhh but no one is answering my questions

OpenStudy (anonymous):

Nice, I think I got the questions right. It is a bit tedious typing them in, thank you for your help, friend.

OpenStudy (aum):

You are welcome.